OpenStudy (anonymous):

Ronny is making 5 batches of lemonade. Each batch requires 2 and 1/3 quarts of water. How much water will Ronny need?

OpenStudy (anonymous):

Please help! I will give a medal to the first helper!

OpenStudy (anonymous):

@brainbuzzed5646 @EthanLaweka anyoneeee

OpenStudy (anonymous):

do 5 times 2 and 1/3

OpenStudy (anonymous):

2 1/3=7/3 7/3*5=35/3 or 11.66666667 quarts of water for 5 batches

OpenStudy (anonymous):

which is is 11 and 2/3

OpenStudy (anonymous):

thank you guys :D

OpenStudy (anonymous):

since I gave Nae Nae the medal, I'll be your fan @cutepochacco :D

OpenStudy (anonymous):

Thanks for the medal!

OpenStudy (anonymous):

You're welcome!